# Learning Encoding-Decoding Direction Pairs to Unveil Concepts of Influence in Deep Vision Networks

### v1

@@ 3 Background
The latent factor of a concept is a scalar linked to the concept’s presence, embedded in the latent space via multiplication with its encoding direction, also called the concept’s signal direction. For this reason, we also refer to this latent factor as the signal value. Feature activations are considered as linear combinations of signals and noisy directional components called distractors. In the proposed approach, a filter is a decoding direction that, through the inner product with a feature representation, extracts the signal’s value. Below we provide a more formal explanation of these terms and provide details essential to understand our contributions.
@@ 3.2 Signals, Distractors, Filters, Concept-Detectors
In encoding a single concept $i$, Kindermans et al. (2017); Pahde et al. (2024) proposed a model for the data generation process of feature representations: ${\bm{x}}_{{\bm{p}}}=\alpha_{{\bm{p}}}{{\bm{s}}}_{i}+\beta_{{\bm{p}}}{\bm{d}},\,{\bm{s}}_{i},{\bm{d}}\in{\mathbb{R}}^{D},\alpha_{{\bm{p}}},\beta_{{\bm{p}}}\in{\mathbb{R}}$. Here, ${{\bm{s}}}_{i}$ is the signal direction that carries the information of whether ${\bm{x}}_{{\bm{p}}}$ is part of concept $i$. The concept information lies within the signal value $\alpha_{{\bm{p}}}$. Larger $\alpha_{{\bm{p}}}$ suggests greater confidence that ${\bm{x}}_{{\bm{p}}}$ belongs to concept $i$. $\bm{d}$} is the distractor direction, modeling noise, or information not related to the concept. $\beta_{{\bm{p}}}$ follows a random distribution, typically the gaussian or uniform distribution, and is independent of whether ${\bm{x}}_{{\bm{p}}}$ belongs to concept $i$. According to Kindermans et al. (2017), the value of the signal $\alpha_{{\bm{p}}}$ can be extracted using a regression filter ${\bm{w}}_{i}$ and the inner product: $z_{{\bm{p}},i}={\bm{w}}_{i}^{T}{\bm{x}}_{{\bm{p}}}=\alpha_{{\bm{p}}}{\bm{w}}_{i}^{T}{{\bm{s}}}_{i}+\beta_{{\bm{p}}}{\bm{w}}_{i}^{T}{\bm{d}}$, if we choose ${\bm{w}}_{i}$ : ${\bm{w}}_{i}\perp{\bm{d}}$, and ${\bm{w}}_{i}^{T}{{\bm{s}}}_{i}=1$. Let $\sigma$ denote the sigmoid function. Since stronger values of $\alpha_{{\bm{p}}}$ indicate more confidence in concept presence, when combined with a threshold $b_{i}\in{\mathbb{R}}$ that can be learned from data, this regression filter can be turned into a concept detector: $y_{{\bm{p}},i}=\sigma(z_{{\bm{p}},i}-b_{i})$, essentially a binary classifier that can answer the question of whether $\bm{p}$} belongs to concept $i$.
With access to the signal value, Haufe et al. (2014); Kindermans et al. (2017) offer a formula to estimate the concept’s signal direction:
@@ 3.3 Unsupervised Interpretable Direction Learning
Figure 3: The core concept of Unsupervised Interpretable Basis Extraction (UIBE Doumanoglou et al. (2023)) is to learn a set of concept detectors, which are essentially binary linear classifiers with learnable filters and biases. These detectors aim to transform feature representations to the soft-binary vector space of concepts in which the newly transformed representations are sparse. In this procedure the input to the method is the network features corresponding to images coming from a concept dataset (without the need for annotations) and the only learnable parameters are the previously mentioned filters and biases themselves. Identifying the concept name behind each detector is done in a post-processing step with a procedure we refer to as Direction Labeling.
Recent research (Doumanoglou et al., 2023) introduced an unsupervised method to identify concepts from the structure of the latent space. Motivated by the directional encoding of concepts, the method partitions the latent space into linear regions, each represented by a hyperplane and a normal vector, forming clusters. Feature activations from an unlabeled concept dataset, possibly activations of images from the network’s training set, are assigned to these clusters. The method learns $\bm{W}$} and $\bm{b}$} of a feature-to-cluster membership function, a mapping to the semantic space, with ${\bm{y}}_{{\bm{p}}}=\sigma({\bm{W}}^{T}{\bm{x}}_{{\bm{p}}}-{\bm{b}})\in[0,1]^{I},{\bm{W}}\in{\mathbb{R}}^{D\times I},{\bm{b}}\in{\mathbb{R}}^{I}$, and $I$ as the cluster count. By softly assigning features to a small number of clusters, the interpretability of the clustering is improved. This is grounded in the idea that an image patch generally holds only a few semantic labels from a larger set, reflecting sparsity in the semantic space. Sparsity in the assignments is achieved using two loss terms: the first is Sparsity Loss ($\mathcal{L}^{s}$), and the second is Maximum Activation Loss ($\mathcal{L}^{ma}$), which ensures binary cluster membership:
@@ 3.4 Direction Labeling
When learning directions in an unsupervised manner, the name of the concept represented by each encoding-decoding direction pair is unknown. In a real-world setting, the concept name could be identified by manually inspecting the samples that (maximally) activate the decoding direction. However, in a benchmark setting, where interpretability needs to be quantified in terms of metrics and without human involvement, it is necessary to automate the assignment of a concept label to each of the direction pairs. We refer to this process as direction labeling and employ Network Dissection Bau et al. (2017) for its implementation. In essence, Network Dissection assigns a semantic label to each of the concept detectors based on their segmentation performance in a dataset with annotated concepts. Although it was originally proposed as a method to assign labels to each of the basis vectors in the natural latent space basis, Network Dissection is capable of assigning labels to any set of directions after basis change. Despite possible biases against unsupervised learning due to annotation limitations, we adopt and expand on this labeling protocol as a best-effort approach to evaluate the interpretability of our concept detectors and the rest of the unsupervised state-of-the-art.
@@ 3.5 Concept Sensitivity Testing
Given the intermediate representation of an image belonging to class $k$ and the direction of concept $c$ in the latent space, RCAV (Pfau et al., 2020) measures the sensitivity of the model to concept $c$ when predicting class $k$. This is accomplished by perturbing the representation towards the direction of the concept with strength $\alpha$, and subsequently comparing the output probability of the network for the same class before and after the perturbation. Subsequently, an overall dataset score in the range [1,1][-1,1] is computed, where zero means inconsistent use of the concept by the model, while extremes indicate consistent and strong positive or negative concept contributions to predict class $k$. A statistical test compares concept sensitivity against sensitivity towards random directions to ensure significance. In our evaluations, we refer to directions of significant influence in cases where the directions meet the criteria of this statistical significance test. Although the initial work used CAVs to test concept sensitivity, a recent study (Pahde et al., 2024) suggests that the appropriate direction for this purpose is the concept’s encoding direction, an aspect that we consider in our evaluations.
4 Method
@@
Figure 4: The proposed method analyzes the latent space to uncover its directional structure. Because many concepts are naturally encoded as specific directions, this process often reveals the encoding-decoding mechanism of meaningful, monosemantic, and highly interpretable concepts. The figure depicts an overview of the method’s components. $\mathcal{L}$ denotes loss terms. Purple indicates contributions of this work, while light gray indicates loss terms from Doumanoglou et al. (2023; 2024).
For a specific network layer, our method receives as input the feature representations of images sourced from a concept dataset, that is, an unlabeled dataset of images coming from the domain of the network depicting concepts that the network might use to make predictions, and utilizes the upper part of the network $f^{+}$ after the layer of study. The aim of our approach is to read the structure of the embedding space in the specific layer and learn encoding-decoding direction pairs that explain this structure. Since concepts are often encoded in latent space directions, reading the structure may reveal the mechanisms to encode and decode meaningful, monosemantic, and highly interpretable concepts. The method is unsupervised, and therefore, the identified direction pairs may correspond to concepts that do not align with human intuition. However, they reflect clear directional clusters in the latent space of the network. Thus, this approach has the potential to reveal erroneous strategies exploited by the model to make predictions (Section 7.11). In an attempt to make these clusters as meaningful as possible, we optimize for a sparsity property of interpretability in the feature-to-cluster assignments (Section 3.3). Additionally, to ensure that those clusters are influential to the network’s predictions, we exploit the uncertainty region of the model.
We extend the previous signal-distractor data model (Section 3.2) from the encoding of a single concept to multiple concepts (Section 4.1) and learn concept detectors $\{{\bm{W}},{\bm{b}}\}$ using the objectives of Section 3.3. In the new data model, we remove the constraints of Doumanoglou et al. (2023) regarding feature space standardization and the orthogonality between filters, allowing for a more flexible clustering which, as we show in the experiments, reduces redundancy. However, when lifting those constraints, we found that they additionally acted as regularizers that prevented direction collapse and trivial clustering; thus, we address their removal with additional loss terms discussed in Section 4.2 that sustain or even improve the interpretability of the clustering. We additionally estimate concept signal directions using learnable signal vectors $\hat{{\bm{s}}}_{i}$ (Section 4.3). Furthermore, we propose Uncertainty Region Alignment (Section 4.4), a loss that can assist in improving the interpretability of the directions or in aligning signal vectors with directions of significant influence. Finally, we propose to learn the direction pairs using $\epsilon$-constrained optimization via the Augmented Lagrangian Loss which we describe in Section 4.5. In summary, concept detectors and signal vectors are learned together in an end-to-end process, influenced by the losses of Sections 3.3, 4.2, 4.3, and 4.4 under the optimization scheme of Section 4.5. An overview of our method and the interconnections between its components is provided in Figures 4 and 5.
@@ 4.1 Multi-Concept Signal-Distractor Data Model
We introduce an extended signal-distractor data model for the latent space, which models the encoding of multiple concepts. Each patch embedding ${\bm{x}}_{{\bm{p}}}$ is considered as a linear combination of latent concept signals ${\bm{S}}\in{\mathbb{R}}^{D\times I}$ and distractors ${\bm{D}}\in{\mathbb{R}}^{D\times F},F\leq D-I$. We also consider a latent space bias ${\bm{c}}\in{\mathbb{R}}^{D}$, common for all ${\bm{x}}_{{\bm{p}}}$.
@@
with $\bm{\alpha}_{{\bm{p}}}\in{\mathbb{R}}^{I}$ and ${\bm{\beta}}_{{\bm{p}}}\in{\mathbb{R}}^{F}$. $\bm{S}$} is a matrix of $I\in{\mathbb{N}}^{+}$, $D$-dimensional, unit-norm concept signal directions and $\bm{D}$} a matrix denoting a basis for distractor components. Each signal direction encodes the presence of a distinct concept. We apply the same assumptions for individual signal values ${\alpha}_{{\bm{p}},i}$ (the $i$-th element of $\bm{\alpha}_{{\bm{p}}}$) and distractor coefficients ${\beta}_{{\bm{p}},f}$ as in Section 3.2. Finally, we further assume that only a limited number of semantic concepts are assigned to ${\bm{x}}_{{\bm{p}}}$, among many possible semantic labels.
Figure 5: Left: The learnable parameters of the method $\hat{{\bm{S}}},{\bm{W}},{\bm{b}}$ and intermediate variables $\mathbf{z}_{{\bm{p}}},{\bm{y}}_{{\bm{p}}},{\bm{q}}_{{\bm{p}}}$. Top Right: Feature manipulation and Uncertainty Region Alignment. Bottom Right: Loss terms $\mathcal{L}$ with their dependencies. Purple indicates loss contributions of this work, while light gray indicates loss terms from Doumanoglou et al. (2023; 2024).
4.2 Interpretability Losses to Recover Implicit Regularizations
@@
Excessively Active Classifier Loss ($\mathcal{L}^{eac}$) This loss penalizes excessively large clusters to prevent trivial solutions where all inputs are assigned to a single cluster. It relies on a hyper-parameter $\rho\in(0,1)$, similar to sparse autoencoders Ng et al. (2011), which sets a proportional bound on cluster size. The unreduced formula for the $i$-th cluster is below, with $\gamma>1,\gamma\in{\mathbb{R}}^{+}$ as a sharpening factor, $1-\rho$ normalizing the loss in the range $[0,1]$ and $\bm{p}$} varying across all pixels and image representations in the concept dataset:
@@
Focal Sparsity Loss ($\mathcal{L}^{fs}$) Inspired by Focal Loss (Lin et al. (2017)), we introduce Focal Sparsity Loss, which puts more emphasis on sparsifying the feature-to-cluster assignments of the most challenging patch embeddings. To this end, for each ${\bm{x}}_{{\bm{p}}}$ we calculate a coefficient $\theta_{{\bm{p}}}\in[0,1]$, which is related to the number of clusters assigned to ${\bm{x}}_{{\bm{p}}}$:
@@
with $\mu$ a sharpening factor and ${\bm{q}}_{{\bm{p}}}\in[0,1]$ an intermediate variable with elements inversely proportional to the patch’s number of cluster assignments, calculated as in (2). If $\mathcal{L}^{s}_{{\bm{p}}}$ (2) denotes the Sparsity Loss for pixel $\bm{p}$}, the Focal Sparsity Loss is defined as:
Similar to the previous case, in (7) $\bm{p}$} varies across all spatial elements and image representations in the concept dataset. When we use Focal Sparsity Loss we use it as a replacement for the Sparsity Loss $\mathcal{L}^{s}$ of Doumanoglou et al. (2023).

### v2

@@ 3 Background
In our formulation, which follows the linear representation hypothesis (Superposition; LinearRepresentationsNLP), the latent factor of a concept is a scalar denoting the concept’s presence in the input, and is embedded in the latent space via multiplication with the concept’s encoding direction, also called the concept’s embedding or signal direction (PatternNet). For this reason, we also refer to this latent factor as the signal value. Feature activations are considered as linear combinations of signals and noisy directional components called distractors. In the proposed approach, a filter, through the inner product with an input representation, can extract the signal’s value and for this reason we term it a decoding direction. Below we provide a more formal explanation of these terms and provide details essential to understand our contributions.
@@ 3.2 Signals, Distractors, Filters, Concept-Detectors
In encoding a single concept $i$, PatternNet; PCAV proposed a model for the data generation process of feature representations: ${\bm{x}}_{{\bm{p}}}=\alpha_{{\bm{p}}}{{\bm{s}}}_{i}+\beta_{{\bm{p}}}{\bm{d}},\,{\bm{s}}_{i},{\bm{d}}\in{\mathbb{R}}^{D},\alpha_{{\bm{p}}},\beta_{{\bm{p}}}\in{\mathbb{R}}$. Here, ${{\bm{s}}}_{i}$ is the signal direction that carries the information of whether ${\bm{x}}_{{\bm{p}}}$ is part of concept $i$. The concept information lies within the signal value $\alpha_{{\bm{p}}}$. Larger $\alpha_{{\bm{p}}}$ suggests greater confidence that ${\bm{x}}_{{\bm{p}}}$ belongs to concept $i$. $\bm{d}$} is the distractor direction, modeling noise, or information not related to the concept. $\beta_{{\bm{p}}}$ follows a random distribution, typically the gaussian or uniform distribution, and is independent of whether ${\bm{x}}_{{\bm{p}}}$ belongs to concept $i$. According to PatternNet, the value of the signal $\alpha_{{\bm{p}}}$ can be extracted using a filter ${\bm{w}}_{i}$ and the inner product: $z_{{\bm{p}},i}={\bm{w}}_{i}^{T}{\bm{x}}_{{\bm{p}}}=\alpha_{{\bm{p}}}{\bm{w}}_{i}^{T}{{\bm{s}}}_{i}+\beta_{{\bm{p}}}{\bm{w}}_{i}^{T}{\bm{d}}$, if we choose ${\bm{w}}_{i}$ : ${\bm{w}}_{i}\perp{\bm{d}}$, and ${\bm{w}}_{i}^{T}{{\bm{s}}}_{i}=1$. Let $\sigma$ denote the sigmoid function. Since stronger values of $\alpha_{{\bm{p}}}$ indicate more confidence in concept presence, when combined with a threshold $b_{i}\in{\mathbb{R}}$ that can be learned from data, this filter can be turned into a concept detector: $y_{{\bm{p}},i}=\sigma(z_{{\bm{p}},i}-b_{i})$, essentially a binary classifier that can answer the question of whether $\bm{p}$} belongs to concept $i$.
With access to the signal value, NeuroImaging; PatternNet offer a formula to estimate the concept’s signal direction:
@@ 3.3 Unsupervised Interpretable Direction Learning
Figure 3: The core concept of Unsupervised Interpretable Basis Extraction (UIBE UIBE) is to learn a set of concept detectors, which are essentially binary linear classifiers with learnable filters and biases. These detectors aim to transform feature representations to the soft-binary vector space of concepts in which the newly transformed representations are sparse. In this procedure the input to the method consists of image representations coming from an unlabeled concept dataset. Identifying the concept name behind each detector is done in a post-processing step with a procedure we refer to as Direction Labeling.
Recent research (UIBE) introduced an unsupervised method to identify concepts from the structure of the latent space. Motivated by the directional encoding of concepts, the method partitions the latent space into linear regions, each represented by a hyperplane and a normal vector, forming clusters. Feature activations from an unlabeled concept dataset, possibly activations of images from the network’s training set, are assigned to these clusters. The method learns $\bm{W}$} and $\bm{b}$} of a feature-to-cluster membership function, a mapping to the semantic space, with ${\bm{y}}_{{\bm{p}}}=\sigma({\bm{W}}^{T}{\bm{x}}_{{\bm{p}}}-{\bm{b}})\in[0,1]^{I},{\bm{W}}\in{\mathbb{R}}^{D\times I},{\bm{b}}\in{\mathbb{R}}^{I}$, and $I$ as the cluster count. By softly assigning features to a small number of clusters, the interpretability of the clustering is improved. This is grounded in the idea that an image patch generally holds only a few semantic labels from a larger set, reflecting sparsity in the semantic space. Sparsity in the assignments is achieved using two loss terms: the first is Sparsity Loss ($\mathcal{L}^{s}$), and the second is Maximum Activation Loss ($\mathcal{L}^{ma}$), which ensures binary cluster membership:
@@ 3.4 Direction Labeling
When learning directions in an unsupervised manner, the name of the concept represented by each encoding-decoding direction pair is unknown. In a real-world setting, the concept name could be identified by manually inspecting the samples that (maximally) activate the decoding direction. However, in a benchmark setting, where interpretability needs to be quantified in terms of metrics and without human involvement, it is necessary to automate the assignment of a concept label to each of the direction pairs. We refer to this process as direction labeling and employ Network Dissection NetworkDissection for its implementation. In essence, Network Dissection assigns a semantic label to each of the concept detectors based on their segmentation performance in a dataset with annotated concepts. Although it was originally proposed as a method to assign labels to each of the basis vectors in the natural latent space basis, Network Dissection is capable of assigning labels to any set of directions after basis change. Despite possible biases against unsupervised learning due to annotation limitations, we adopt and expand on this labeling protocol as a best-effort approach to evaluate the interpretability of our concept detectors and the rest of the unsupervised baselines.
@@ 3.5 Concept Sensitivity Testing
Given the intermediate representation of an image belonging to class $k$ and the direction of concept $c$ in the latent space, RCAV (RCAV) measures the sensitivity of the model to concept $c$ when predicting class $k$. This is accomplished by perturbing the representation towards the direction of the concept with strength $\alpha$, and subsequently comparing the output probability of the network for the same class before and after the perturbation. Subsequently, an overall dataset score in the range [1,1][-1,1] is computed, where zero means inconsistent use of the concept by the model, while extremes indicate consistent positive or negative concept contributions to predicting class $k$. A statistical test compares concept sensitivity against sensitivity towards random directions to ensure significance. In our evaluations, we refer to directions of significant influence in cases where the directions meet the criteria of this statistical significance test. Although the initial work used CAVs to test concept sensitivity, a recent study (PCAV) suggests that the appropriate direction for this purpose is the concept’s encoding direction, an aspect that we consider in our evaluations.
4 Proposed Method
@@
Figure 4: The proposed method analyzes the latent space to uncover its directional structure. Because many concepts are naturally encoded as specific directions, this process often reveals the encoding-decoding mechanism of meaningful, monosemantic, and highly interpretable concepts. The figure depicts an overview of the method’s components. $\mathcal{L}$ denotes loss terms. Purple indicates contributions of this work, while light gray indicates loss terms from UIBE; CBE.
Method Overview
The aim of our unsupervised approach is to reveal the encoding-decoding mechanism of meaningful concepts that influence the network’s predictions. Our method learns encoding-decoding direction pairs that explain the structure of a deep vision network’s latent space and takes as input feature representations from an unlabeled concept dataset. The latent space, from which feature representations come, corresponds to a network layer at a given depth. We jointly learn the decoding directions (concept detectors) $\{{\bm{W}},{\bm{b}}\}$ and the encoding directions (signal vectors) $\hat{{\bm{S}}}$ in an end-to-end process.
Our formulation is built upon several key contributions:
  1. 1.
    Flexible and Sparser Clustering: We lift previous constraints on the concept detectors (such as orthogonality between filters and feature space standardization) in UIBE, enabling more flexible and less redundant (i.e., sparser) clustering. To preserve or even improve interpretability, we incorporate two novel loss terms: Focal Sparsity Loss and Excessively Active Classifier Loss (Section 4.2).
  2. 2.
    Encoding Direction Estimation: We incorporate an extended multi-concept signal-distractor model (Section 4.1) and define signal vectors ($\hat{\mathbf{S}}$) as estimators for the concept encoding directions based on a probabilistic framework (Section 4.3).
  3. 3.
    Network Instruction Driven Search: We introduce Uncertainty Region Alignment ($\mathcal{L}^{ur}$) (Section 4.4), a novel loss that can improve concept interpretability and influence by aligning concept detector ambiguity with model prediction uncertainty. For the latter, the method makes use of the upper part of the network, after the layer of study, that produces output class probabilities, denoted as $f^{+}$.
  4. 4.
    Effective Optimization: The entire set of learning parameters is optimized jointly using an $\epsilon$-constrained optimization scheme based on the Augmented Lagrangian Loss (Section 4.5), which allows direct control of clustering quality with easier to tune, interpretable hyper-parameters.
Our method overview and component interconnections are shown in Figures 4 and 5. The used loss terms are described in Table 1, and the learning algorithm is outlined in Section A.2
@@ 4.1 Multi-Concept Signal-Distractor Data Model
We introduce an extended signal-distractor data model for the latent space, which models the encoding of multiple concepts. Each patch embedding ${\bm{x}}_{{\bm{p}}}$ is considered as a linear combination of latent concept signals ${\bm{S}}\in{\mathbb{R}}^{D\times I}$ and distractors ${\bm{D}}\in{\mathbb{R}}^{D\times F}$. We also consider a latent space bias ${\bm{c}}\in{\mathbb{R}}^{D}$, common for all ${\bm{x}}_{{\bm{p}}}$.
@@
with $\bm{\alpha}_{{\bm{p}}}\in{\mathbb{R}}^{I}$ and ${\bm{\beta}}_{{\bm{p}}}\in{\mathbb{R}}^{F}$. $\bm{S}$} is a matrix of $I\in{\mathbb{N}}^{+}$, $D$-dimensional, unit-norm concept signal directions and $\bm{D}$} a matrix denoting a basis for distractor components. Each signal direction encodes the presence of a distinct concept. We apply the same assumptions for individual signal values ${\alpha}_{{\bm{p}},i}$ (the $i$-th element of $\bm{\alpha}_{{\bm{p}}}$) and distractor coefficients ${\beta}_{{\bm{p}},f}$ as in Section 3.2.
Figure 5: Left: The learnable parameters of the method $\hat{{\bm{S}}},{\bm{W}},{\bm{b}}$ and intermediate variables $\mathbf{z}_{{\bm{p}}},{\bm{y}}_{{\bm{p}}},{\bm{q}}_{{\bm{p}}}$. Top Right: Feature manipulation and Uncertainty Region Alignment. Bottom Right: Loss terms $\mathcal{L}$ with their dependencies. Purple indicates loss contributions of this work, while light gray indicates loss terms from UIBE; CBE.
4.2 Losses to Preserve or Improve Interpretability
@@
Excessively Active Classifier Loss ($\mathcal{L}^{eac}$) This loss penalizes excessively large clusters to prevent trivial solutions in which all inputs are assigned to a single cluster. It relies on a hyper-parameter $\rho\in(0,1)$, similar to sparse autoencoders sparse-autoencoder, which sets a proportional bound on cluster size. With ${y}_{{\bm{p}},i}\in[0,1]$ denoting the membership of patch $\bm{p}$} for cluster $i$ (Section 3.3), the unreduced loss formula for the $i$-th cluster is below, with $\gamma>1,\gamma\in{\mathbb{R}}^{+}$ as a sharpening factor, $1-\rho$ normalizing the loss in the range $[0,1]$, and $\bm{p}$} varying across all pixels and image representations in the concept dataset:
@@
Focal Sparsity Loss ($\mathcal{L}^{fs}$) Inspired by Focal Loss (focal-loss), we introduce Focal Sparsity Loss, which places more emphasis on sparsifying the feature-to-cluster assignments of the most challenging patch embeddings. If $\mathcal{L}^{s}_{{\bm{p}}}$ (2) denotes the Sparsity Loss for pixel $\bm{p}$}, the Focal Sparsity Loss is defined as:
i.e. based on a weighted scheme defined by coefficients $\theta_{{\bm{p}}}\in[0,1]$. We choose $\theta_{{\bm{p}}}$ to be proportional to the patch’s number of cluster assignments. In detail:
@@
[…]